OpenStudy (anonymous):

The base and height of a triangle sum up to 15, and the area of the triangle is 24. What are the dimensions of the triangle?

OpenStudy (anonymous):

you mean the base + height = 15 ?

OpenStudy (anonymous):

Yes!

OpenStudy (anonymous):

Okay,you just need to create a system!

OpenStudy (anonymous):

\(\sf {The \ Base \Longrightarrow \ x} \) \(\sf {The \ Height \Longrightarrow \ y} \)

OpenStudy (anonymous):

and this is the system:

OpenStudy (anonymous):

\(\sf \large x + y = 15\) \(\sf \large\frac{ xy }{ 2 } = 24\)

OpenStudy (anonymous):

understood?

OpenStudy (anonymous):

Yes, thank you!

OpenStudy (anonymous):

and now can you solve that system madam?

OpenStudy (anonymous):

Yes, I just use the process of substitution correct?

OpenStudy (anonymous):

yes,it will work :)

OpenStudy (anonymous):

Okay, I will try as well! :)

OpenStudy (anonymous):

nice;)

OpenStudy (anonymous):

I got a crazy number :o

OpenStudy (anonymous):

need help?

OpenStudy (anonymous):

I got y=15-x and tried to substitute that into the other equation and just made a mess!

OpenStudy (anonymous):

me too :P are you sure that the question is not wrong?

OpenStudy (anonymous):

it seems that it's a quadratic equation

OpenStudy (anonymous):

i'm not sure about the question! how did you find this question?

OpenStudy (anonymous):

It's the right equation :/

OpenStudy (anonymous):

An example I was given

OpenStudy (anonymous):

@Anecia23

OpenStudy (anonymous):

@ganeshie8 , what's your idea about it?

OpenStudy (anonymous):

@hartnn

hartnn (hartnn):

yes, it will be quadratic product = 12 sum =15 x^2 + 15x+12 =0

OpenStudy (anonymous):

but it will be \(\large \sqrt{177}\)

hartnn (hartnn):

eh, in terms of sqrt 177.... the dimensions are not integers....but irrational numbers..

OpenStudy (anonymous):

so don't you think that the question is wrong?

hartnn (hartnn):

if the question just belongs to quadratic equations, then no harm in dimensions being irrational...

OpenStudy (nirmalnema):

ohh..ok

OpenStudy (nirmalnema):

let the height be (x) so base is (15-x) area = 1/2*base*height 24=1/2 (x)(15-x) x^2-15x+48=0

OpenStudy (anonymous):

and this time,it will be \(\large \sqrt{33}\)

hartnn (hartnn):

yeah, still irrational

OpenStudy (nirmalnema):

@hartnn how do u get 12

hartnn (hartnn):

its 48 only my bad

ganeshie8 (ganeshie8):

if u dont want a quadratic, \( x+y = 15\) -------(1) \(\frac{2}{2} xy = 24\) \(x-y = \sqrt{15^2-4 \times 48}\) \(x-y = \sqrt{33}\) ---(2)

OpenStudy (nirmalnema):

ohh///.....okk

hartnn (hartnn):

effect of doing thing in head.... :P

hartnn (hartnn):

xy should be 48....

OpenStudy (nirmalnema):

hehe......no problem same with me.....i missedd half

OpenStudy (anonymous):

guys it's not important :D the important thing is "the question is irrational"

ganeshie8 (ganeshie8):

wats problem wid irrational lengths ?

OpenStudy (nirmalnema):

so there is no problem in it..

ganeshie8 (ganeshie8):

my height is irrational its 174.7834232398123442223... cm

hartnn (hartnn):

lol

OpenStudy (anonymous):

XD

hartnn (hartnn):

how do you have so perfect instrument to measure it ? :P

OpenStudy (anonymous):

but @ganeshie8 usually , teachers give "rational questions"

hartnn (hartnn):

lets ask @Anecia23 what she got ?

OpenStudy (anonymous):

or when they give "irrational questions" they will say "you can't find the exact value of ... "

hartnn (hartnn):

how do you have so perfect instrument to measure it ? :P

ganeshie8 (ganeshie8):

ive put myself vertically before a scanning electron miscroscope... that gave my precision of 1nm... after that did a simple extrapolation to 23 sigfigs lol

hartnn (hartnn):

seems legit

OpenStudy (nirmalnema):

u got great height........=P

OpenStudy (anonymous):

why is there so many people on one question

OpenStudy (anonymous):

i don't know too :D sometimes this will happen :D

hartnn (hartnn):

because beautiful profile pics attract more helpers.. :P

OpenStudy (anonymous):

this an good idea too XD

OpenStudy (nirmalnema):

haha....

hartnn (hartnn):

i came here for that reason only.....not because i was called :P

OpenStudy (anonymous):

so i won't call you again :D

hartnn (hartnn):

call me only when the profic pic is of gorgeous girl ;)

OpenStudy (anonymous):

:D

ganeshie8 (ganeshie8):

lol here is my final attempt to make the gorgeous girl happy :) the complete solution \( x+y = 15\) -------(1) \(\frac{1}{2} xy = 24\) \(x-y = \sqrt{15^2-4 \times 48}\) \(x-y = \sqrt{33}\) ---(2) (1) + (2) \(2x = 15 + \sqrt{33}\) \(\color{red}{x = \frac{15 + \sqrt{33}}{2}}\) (1) - (2) \(2y = 15 - \sqrt{33}\) \(\color{red}{y = \frac{15 - \sqrt{33}}{2}}\) if the question asks for exact value, then @PFEH.1999 , as u said we must lev the answer in radical form as above. otherwise we can plugit in calculator and approximate

OpenStudy (anonymous):

nice job :)

hartnn (hartnn):

where's the gorgeous girl now? does she even know whats going on here :P

OpenStudy (anonymous):

@Anecia23 , got it?

OpenStudy (anonymous):

so you can't find the exact value of x

OpenStudy (anonymous):

and y

hartnn (hartnn):

i would take a moment to welcome her here ... \[ \begin{array}l\color{red}{\text{W}}\color{orange}{\text{E}}\color{#e6e600}{\text{L}}\color{green}{\text{C}}\color{blue}{\text{O}}\color{purple}{\text{M}}\color{purple}{\text{E}}\color{red}{\text{ }}\color{orange}{\text{t}}\color{#e6e600}{\text{o}}\color{green}{\text{ }}\color{blue}{\text{O}}\color{purple}{\text{P}}\color{purple}{\text{E}}\color{red}{\text{N}}\color{orange}{\text{S}}\color{#e6e600}{\text{T}}\color{green}{\text{U}}\color{blue}{\text{D}}\color{purple}{\text{Y}}\color{purple}{\text{!}}\color{red}{\text{!}}\color{orange}{\text{}}\end{array} \ddot \smile \]

OpenStudy (anonymous):

Sorry guys I fell asleep! :o

hartnn (hartnn):

she's here! she's here :D

OpenStudy (anonymous):

I appreciate the help! Thank you so much to all of you!

OpenStudy (anonymous):

You're welcome :)

OpenStudy (anonymous):

i forgot to welcome her to openstudy too :D besides that i forgot my the welcome code... XD

OpenStudy (anonymous):

Hehe! First time using this website! I was having problems with this question but it makes sense now! Thanks again to everybody how helped! :*

OpenStudy (anonymous):

:);)

hartnn (hartnn):

hey, if you have any doubts browsing this site, you can ask me :) *us

OpenStudy (anonymous):

or post it in this group: http://openstudy.com/study#/groups/openstudy%20feedback :)

OpenStudy (anonymous):

I'll definitely be back soon! :D

hartnn (hartnn):

yaaay :)

OpenStudy (anonymous):

:)

OpenStudy (anonymous):

^___^

OpenStudy (anonymous):

Any question ???

OpenStudy (anonymous):

oh! @E.ali did you see last comments? there's not any question :)

OpenStudy (anonymous):

hehehe !